# Large Deviations Principle for Isoperimetry and Its Equivalence to Nonlinear Log-Sobolev Inequalities

### v3

@@ 3.1. From Transport to Variant Isoperimetry
$\nu_{X}^{\otimes n^{\prime\prime}}(B^{c})$ $=\mathbb{P}\{\sum_{i=n^{\prime}+1}^{n}d^{p}(X_{i},x_{0})>r^{p}\}\leq\mathbb{E}_{\nu}[e^{\lambda d^{p}(X,x_{0})}]^{n^{\prime\prime}}e^{-\lambda r^{p}}=e^{-\lambda\left(r^{p}-n^{\prime\prime}\ell_{\lambda}\right)},$
where $\ell_{\lambda}:=\frac{1}{\lambda}\ln\mathbb{E}_{\nu}[e^{\lambda d^{p}(X,x_{0})}]$ is finite for all $\lambda>0$ by assumption.
Choose $r=(n\delta^{\prime})^{1/p}$. We then have
(3.31) $\nu_{X}^{\otimes n}(B^{c})$ $\leq e^{-\lambda n\left(\delta^{\prime}-\epsilon^{\prime}\ell_{\lambda}\right)}.$
For fixed $\delta^{\prime}>0$ and $\beta>\alpha^{\prime}$, we choose $\lambda$ large enough and $\epsilon^{\prime}$ small enough so that
(3.32) $\lambda\left(\delta^{\prime}-\epsilon^{\prime}\ell_{\lambda}\right)$ $>\beta\textrm{ and }\mathbb{E}_{\nu}[d^{p}(Y,x_{0})]<\frac{\delta^{\prime}}{\epsilon^{\prime}}.$
By the law of large numbers, $\nu_{Y}^{\otimes n^{\prime\prime}}(B)\to 1$ as $n\to\infty$.
For $\mathbf{u}\in\mathcal{X}^{n^{\prime}}$, denote $\tilde{A}_{\mathbf{u}}:=\{\mathbf{v}\in\mathcal{X}^{n^{\prime\prime}}:(\mathbf{u},\mathbf{v})\in\tilde{A}\}$, which is the projection of $\tilde{A}$ onto the last $n^{\prime\prime}$ components. Denote $\hat{A}:=\{\mathbf{u}\in\mathcal{X}^{n^{\prime}}:\nu_{X}^{\otimes n^{\prime\prime}}(\tilde{A}_{\mathbf{u}})\geq e^{-n\beta}\}\subseteq F^{n^{\prime}}$. Denote $\bar{A}:=\{(\mathbf{u},\mathbf{v})\in\tilde{A}:\mathbf{u}\in\hat{A}\}$. We then have
[…]
which implies
(3.33) $\nu_{X}^{\otimes n}(\bar{A})\geq e^{-n\alpha^{\prime}}-e^{-n\beta}=e^{-n(\alpha^{\prime}+o(1))},$
with $o(1)$ denoting a term that vanishes as $n\to\infty$.
From (3.31) and (3.32), we see that $B$ and $\tilde{A}_{\mathbf{u}}$ have nonempty intersection for each $\mathbf{u}\in\hat{A}$. Hence, $(\tilde{A}_{\mathbf{u}})^{2r}:=\{\mathbf{v}^{\prime}\in\mathcal{X}^{n^{\prime\prime}}:d(\mathbf{v}^{\prime},\mathbf{v})\leq 2r,\mathbf{v}\in\tilde{A}_{\mathbf{u}}\}$ covers the whole $B$, where $r$ is the radius in (3.30) and set to $(n\delta^{\prime})^{1/p}$. So,
$A^{c_{n}\leq t}\supseteq\tilde{A}^{c_{n}\leq t}$ $=\bigcup_{\mathbf{x}\in\tilde{A}}\{\mathbf{x}\}^{c_{n}\leq t}\supseteq\bigcup_{\mathbf{u}\in\hat{A}}\Big(\{\mathbf{u}\}^{c_{n^{\prime}}\leq t-2^{p}n\delta^{\prime}}\times A_{\mathbf{u}}^{c_{n^{\prime\prime}}\leq 2^{p}n\delta^{\prime}}\Big)$
$\supseteq\bigcup_{\mathbf{u}\in\hat{A}}\Big(\{\mathbf{u}\}^{c_{n^{\prime}}\leq t-2^{p}n\delta^{\prime}}\times B\Big)=\hat{A}^{c_{n^{\prime}}\leq t-2^{p}n\delta^{\prime}}\times B.$
Summarizing the above, it holds that
[…]
$\nu_{Y}^{\otimes n}(A^{c_{n}\leq t})$ $\geq\nu_{Y}^{\otimes n^{\prime}}(\hat{A}^{c_{n^{\prime}}\leq t-2^{p}n\delta^{\prime}})\nu_{Y}^{\otimes n^{\prime\prime}}(B)\geq(1-o(1))\nu_{Y}^{\otimes n^{\prime}}(\hat{A}^{c_{n^{\prime}}\leq t-2^{p}n\delta^{\prime}}),$
where $o(1)$ are terms that vanish as $n\to\infty$. Setting $t=n\tau$, we then have that
$\nu_{Y}^{\otimes n}(A^{n\tau})$ $\geq(1-o(1))\inf_{\hat{A}\subseteq F^{n^{\prime}}:\nu_{X}^{\otimes n^{\prime}}(\hat{A})\geq e^{-n^{\prime}\alpha^{\prime\prime}}}\nu_{Y}^{\otimes n^{\prime}}(\tilde{A}^{c_{n^{\prime}}\leq n^{\prime}\tau^{\prime}}),$
where $\tau^{\prime}:=\frac{\tau-2^{p}\delta^{\prime}}{1-\epsilon^{\prime}}$ and $\alpha^{\prime\prime}:=\frac{\alpha+\delta+H(\epsilon^{\prime})+o(1)}{1-\epsilon^{\prime}}$. That is,
$E_{n}(\alpha,\tau|\nu_{X},\nu_{Y})$ $\leq E_{n^{\prime}}(\alpha^{\prime\prime}+\ln\nu_{X}(F),\tau^{\prime}|\nu_{X|F},\nu_{Y})+o(1),$
where $\nu_{X|F}:=\nu_{X|X\in F}:=\nu_{X}(\cdot|F)$ is the conditional probability measures of $\nu_{X}$ on $F$.
Since $F$ is compact, applying the upper bound on the isoperimetric exponent in Theorem 9, we obtain that
(3.34) $\limsup_{n\to\infty}E_{n}(\alpha,\tau)$ $\leq\limsup_{n\to\infty}E_{n^{\prime}}(\alpha^{\prime\prime}+\ln\nu_{X}(F),\tau^{\prime}|\nu_{X|F},\nu_{Y})$
$\leq\psi(\alpha^{\prime\prime}+\ln\nu_{X}(F),\tau^{\prime}|\nu_{X|F},\nu_{Y})\leq\psi(\alpha^{\prime\prime},\tau^{\prime}|\nu_{X},\nu_{Y}).$
Letting $\epsilon^{\prime}\to 0,\delta\to 0$, and $\delta^{\prime}\to 0$, (i.e., $\alpha^{\prime\prime}\to\alpha$ and $\tau^{\prime}\to\tau$), we further obtain that
$\limsup_{n\to\infty}E_{n}(\alpha,\tau)$ $\leq\lim_{\alpha^{\prime\prime}\downarrow\alpha,\tau^{\prime}\uparrow\tau}\psi(\alpha^{\prime\prime},\tau^{\prime})=\lim_{\alpha^{\prime\prime}\downarrow\alpha}\psi(\alpha^{\prime\prime},\tau),$
where the equality follows since $\psi$ is convex and hence continuous in $\tau>0$.
We now need the following lemma whose proof is provided later.
Lemma 1.
For $\tau>0$, $\alpha\mapsto\psi(\alpha,\tau)$ is right continuous.
By this lemma, we obtain the desired bound $\limsup_{n\to\infty}E_{n}(\alpha,\tau)\leq\psi(\alpha,\tau)$. ∎
We now prove Lemma 1.
Proof of Lemma 1.
Since $\lim_{\alpha^{\prime}\downarrow\alpha}\psi(\alpha^{\prime},\tau)\geq\psi(\alpha,\tau)$, it suffices to prove $\lim_{\alpha^{\prime}\downarrow\alpha}\psi(\alpha^{\prime},\tau)\leq\psi(\alpha,\tau)$. We now use the standard lower semi-continuity and compactness arguments to prove this point.
Denote $\mathsf{C}=\mathsf{W}_{p}^{p}$. Denote
$g(\pi_{XW},\tau):=\min_{\pi_{Y|W}:\mathsf{C}(\pi_{X|W},\pi_{Y|W}|\pi_{W})\leq\tau}D(\pi_{Y|W}\|\nu_{Y}|\pi_{W}),$
which is convex since both the relative entropy and the optimal transport cost functional are convex. The minimum is attained if $g(\pi_{XW},\tau)<\infty$, since we minimize a lower semi-continuous function over a (weakly) closed set. Notice that $\psi(\alpha,\tau)$ is defined as a supremum of $g(\pi_{XW},\tau)$. So, $\tau\mapsto\psi(\alpha,\tau)$ is convex for $\tau\geq 0$, and hence continuous for $\tau>0$ and $\psi(\alpha,\tau)<\infty$.
To prove that $\alpha\mapsto\psi(\alpha,\tau)$ is continuous for $\tau\geq 0$, it suffices to prove that $\psi(\alpha_{k},\tau)\to\psi(\alpha,\tau)$ for all sequences $\alpha_{k}$ such that $\alpha_{k}\downarrow\alpha$. Denote $\pi_{XW}^{(k)}$ as a distribution approximately attaining the supremization in $\psi(\alpha_{k},\tau)$ within a gap $1/k$. Denote $\pi_{Y|W}^{(k)}$ as a conditional distribution attaining $g(\pi_{XW}^{(k)},\tau)$, i.e., the infimization in $\psi(\alpha_{k},\tau)$. Since $\pi_{XW}^{(k)}$ is in $\{\pi_{XW}:D(\pi_{X|W}\|\nu_{X}|\pi_{W})\leq\alpha\}$, a weakly compact set (e.g., [25, Lemma 6.2.12]), by passing a subsequence, we can assume that $\pi_{XW}^{(k)}$ converges to some distribution $\pi_{XW}^{*}$. Denote $\pi_{Y|W}^{*}$ as a conditional distribution attaining $g(\pi_{XW}^{*},\tau-\delta)$. Since by considering the feasible solution $\pi_{X|W}$ for the minimization in $g(\pi_{XW},\tau)$, one can observe that $g(\pi_{XW},\tau)$ is bounded by the finite number $\alpha$. So, without loss of generality, we assume that $D(\pi_{Y|W=w}^{*}\|\nu)$ is finite for both $w=0$ and $1$.
By the triangle inequality,
$\mathsf{W}_{p}(\pi_{X|W}^{(k)},\pi_{Y|W}^{*}|\pi_{W}^{(k)})$ $\leq\mathsf{W}_{p}(\pi_{X|W}^{(k)},\pi_{X|W}^{*}|\pi_{W}^{(k)})+\mathsf{W}_{p}(\pi_{Y|W}^{*},\pi_{X|W}^{*}|\pi_{W}^{(k)}).$
Since the Wasserstein metric induced by a bounded metric $d$ is compatible with the weak topology, it holds that as $k\to\infty$, $\mathsf{W}_{p}(\pi_{X|W=w}^{(k)},\pi_{X|W=w}^{*})\to 0$ for $w\in{\mathrm{supp}}(\pi_{W}^{*}):=\{w\in\{0,1\}:\pi_{W}^{*}(w)>0\}$. Moreover, $\mathsf{W}_{p}$ is bounded, and hence, $\mathsf{W}_{p}(\pi_{X|W}^{(k)},\pi_{X|W}^{*}|\pi_{W}^{(k)})\to\mathsf{W}_{p}(\pi_{X|W}^{*},\pi_{X|W}^{*}|\pi_{W}^{*})=0$ and $\mathsf{W}_{p}(\pi_{Y|W}^{*},\pi_{X|W}^{*}|\pi_{W}^{(k)})\to\mathsf{W}_{p}(\pi_{Y|W}^{*},\pi_{X|W}^{*}|\pi_{W}^{*})\leq\tau-\delta$, as $k\to\infty$. So, for sufficiently large $k$,
$\mathsf{W}_{p}(\pi_{X|W}^{(k)},\pi_{Y|W}^{*}|\pi_{W}^{(k)})$ $\leq\tau.$
That is, $\pi_{Y|W}^{*}$ is a feasible solution to $g(\pi_{XW}^{(k)},\tau)$, which implies $\psi(\alpha_{k},\tau)-\frac{1}{k}\leq g(\pi_{XW}^{(k)},\tau)\leq D(\pi_{Y|W}^{*}\|\nu_{Y}|\pi_{W}^{(k)})$. We in turn obtain that
@@
(3.35) $\limsup_{k\to\infty}\psi(\alpha_{k},\tau)\leq\limsup_{k\to\infty}D(\pi_{Y|W}^{*}\|\nu_{Y}|\pi_{W}^{(k)})=D(\pi_{Y|W}^{*}\|\nu_{Y}|\pi_{W}^{*})=g(\pi_{XW}^{*},\tau-\delta).$
On the other hand, by the lower semi-continuity of the relative entropy,
$D(\pi_{X|W}^{*}\|\nu_{X}|\pi_{W}^{*})\leq\limsup_{k\to\infty}D(\pi_{X|W}^{(k)}\|\nu_{X}|\pi_{W}^{(k)})\leq\alpha.$
Hence, $\pi_{XW}^{*}$ is a feasible solution to the following supremization:
$\psi(\alpha,\tau-\delta)=\sup_{\pi_{XW}:D(\pi_{X|W}\|\nu_{X}|\pi_{W})\leq\alpha}g(\pi_{XW},\tau-\delta).$
So, $g(\pi_{XW}^{*},\tau-\delta)\leq\psi(\alpha,\tau-\delta)$, which, combined with (3.35), implies that $\limsup_{k\to\infty}\psi(\alpha_{k},\tau)\leq\psi(\alpha,\tau-\delta).$ Letting $\delta\downarrow 0$ and by the continuity of $\tau\mapsto\psi(\alpha,\tau)$, it holds that $\limsup_{k\to\infty}\psi(\alpha_{k},\tau)\leq\psi(\alpha,\tau)$. ∎

### v4

@@ 3.5.4. Step 4: Heat Flow
$\leq\sqrt{\mathbb{E}_{\mathbf{X}^{\prime}\sim\mu}[\mathsf{W}_{2}^{2}\left(\mu_{\mathbf{X}_{t}|\mathbf{X}=\mathbf{X}^{\prime}},\delta_{\mathbf{x}}\right)]}+\sqrt{n\delta}$
$=\sqrt{\mathbb{E}_{\mathbf{X}\sim\mu}\sum_{i=1}^{n_{1}}\mathbb{E}_{\mu}[d^{2}(x_{i},X_{i,t})|X_{i}]}+\sqrt{n\delta}$
$=\sqrt{\sum_{i=1}^{n_{1}}\mathbb{E}_{\mu}[d^{2}(x_{i},X_{i,t})]}+\sqrt{n\delta}$
$=\mathsf{W}_{2}\left(\delta_{\mathbf{x}},\mu_{\mathbf{X}_{t}}\right)+\sqrt{n\delta}.$
Therefore,
$\sqrt{\sum_{i=1}^{n_{1}}\mathbb{E}_{\mu}[d^{2}(x_{i},X_{i,t})|x_{i}]}$ $=\mathsf{W}_{2}\left(\mu_{\mathbf{X}_{t}|\mathbf{X}=\mathbf{x}},\delta_{\mathbf{x}}\right)\leq\mathsf{W}_{2}\left(\delta_{\mathbf{x}},\mu_{\mathbf{X}_{t}}\right)+\sqrt{n\delta}$
$\leq\mathsf{W}_{2}\left(\mu_{\mathbf{X}},\mu_{\mathbf{X}_{t}}\right)+2\sqrt{n\delta},$
where the last inequality follows by the triangle inequality again. Moreover, $\mathbb{E}_{\pi^{*}}[d^{2}(X,Y)]=\mathsf{W}_{2}^{2}(\pi_{X|W},\pi_{X_{t}|W}|\pi_{W})$. Therefore, combining these with $\mathsf{W}_{2}(t)\leq\sqrt{n\tau^{\prime\prime}}$ yields
$n_{0}\mathbb{E}_{\pi}[d^{2}(X,Y)]+\sum_{i=1}^{n_{1}}\mathbb{E}_{\mu}[d^{2}(x_{i},Y)|x_{i}]\leq n(\tau^{\prime\prime}+4\delta+4\sqrt{\delta\tau})\;\mu_{\mathbf{X}}\textrm{-a.e.}$
Choosing $\tau^{\prime\prime}=\tau-9\delta-8\sqrt{\delta\tau}$ yields the expectation requirement in (3.23).
We next focus on the variance requirement. Under $\mathrm{CD}(0,\infty)$, the gradient flow (heat flow) $\rho_{t}$ (given in (2.4)) of Cheeger’s energy on the Hilbert space $L^{2}(\mathcal{X},\nu)$ is identical to the gradient flow $\mu_{t}$ of relative entropy on the Wasserstein metric space in the sense that $\mu_{t}=\rho_{t}\nu$ when $\mu_{0}=\rho_{0}\nu$ [6, Theorem 9.3]. In this case, denote $P_{t}$ as the Markov (heat) semigroup operator given by $P_{t}:\rho_{0}\mapsto\rho_{t}$. If in addition, the space satisfies $\mathrm{RCD}(0,\infty)$, then these two flows satisfy the equation $\frac{\mathrm{d}}{\mathrm{d}t}\rho_{t}=\Delta_{\mathcal{E}}\rho_{t}$ for all $t>0$ where $\Delta_{\mathcal{E}}$ is the infinitesimal generator of $P_{t}$ and $\rho_{t}=P_{t}\rho_{0}$ [7, Theorem 2.13 and Section 6.1], and moreover, in this case, Bakry-Émery curvature-dimension condition $\mathrm{BE}(0,\infty)$ is satisfied by the semigroup $P_{t}$ [8, Theorem 4.17].
By the local Poincaré inequalities in [8, Corollary 2.3, Definition 2.4, Theorem 4.17][10, Theorem 4.7.2], under $\mathrm{RCD}(0,\infty)$ (equivalently, $\mathrm{BE}(0,\infty)$), it holds that for every $f\in W^{1,2}(\mathcal{X},d,\nu)$ with the Sobolev space $W^{1,2}(\mathcal{X},d,\nu)$ defined around (2.3) and for $t>0$,
$P_{t}(f^{2})-(P_{t}f)^{2}\leq 2tP_{t}(|\nabla f|_{*}^{2}),\;\nu\textrm{-a.e. in }\mathcal{X},$
where $|\nabla f|_{*}$ is the minimal weak gradient defined below (2.3) and $f\mapsto|\nabla f|_{*}^{2}$ forms the Carré du champ $\Gamma$ induced by the Dirichlet form $\mathcal{E}=2\mathrm{Ch}$. That is, for every $f\in W^{1,2}(\mathcal{X},d,\nu)$,
(3.28) $\mathrm{Var}_{\mu}(f(X_{t})|X=x)\leq 2t\mathbb{E}_{\mu}[|\nabla f|_{*}^{2}(X_{t})|X=x],\;\nu\textrm{-a.e. }x\in\mathcal{X}.$
Let $f(y)=d^{2}(x,y)$ and for $N>0$, let $f_{N}(y)=\min\{f(y),N^{2}\}=\left(\min\{d(x,y),N\}\right)^{2}$. Then, $|\nabla f_{N}|\leq 2\min\{d(x,y),N\}$ and hence $f_{N}\in W^{1,2}(\mathcal{X},d,\nu)$. Substituting this function into (3.28) and noting that $|\nabla f_{N}|_{*}\leq|\nabla f_{N}|\leq 2d(x,y)$ $\nu$-a.e. $y$, we obtain that
$\mathrm{Var}_{\mu}(f_{N}(X_{t})|X=x)\leq 8t\mathbb{E}_{\mu}[d^{2}(x,X_{t})|X=x],\;\nu\textrm{-a.e. }x\in\mathcal{X}.$
Letting $N\to\infty$ and by the monotone convergence theorem, for $\nu\textrm{-a.e. }x\in\mathcal{X}$,
(3.29) $\mathrm{Var}_{\mu}(d^{2}(x,X_{t})|X=x)\leq 8t\mathbb{E}_{\mu}[d^{2}(x,X_{t})|X=x].$
Using the triangle inequality, for any $x_{0},x\in F$,
$\sqrt{\mathbb{E}_{\mu}[d^{2}(x_{0},X_{t})|X=x]}$ $=\mathsf{W}_{2}(P_{t}\delta_{x},\delta_{x_{0}})$
$\leq\mathsf{W}_{2}(P_{t}\delta_{x},\nu)+\mathsf{W}_{2}(\nu,\delta_{x_{0}})$
$\leq\mathsf{W}_{2}(\delta_{x},\nu)+\mathsf{W}_{2}(\nu,\delta_{x_{0}})$
$\leq\mathsf{W}_{2}(\delta_{x},\delta_{x_{0}})+2\mathsf{W}_{2}(\delta_{x_{0}},\nu)$
(3.30) $\leq\mathrm{diam}(F)+2\mathsf{W}_{2}(\delta_{x_{0}},\nu)=:C,$
where the second inequality follows by $\mathsf{W}_{2}(P_{t}\delta_{x},\nu)\leq\mathsf{W}_{2}(\delta_{x},\nu)$, contraction of Wasserstein metric in [7, (6.1)]. Noting that $C<\infty$, $\mathbb{E}_{\mu}[d^{2}(x_{0},X_{t})|X=x]$ is upper bounded by the finite number $C$ uniformly for all $x\in F$. We then obtain from (3.29) that
(3.31) $\mathbb{E}_{\mu}[d^{4}(x_{0},X_{t})|X=x]\leq C^{4}+8tC^{2}.$
Moreover, by triangle inequality and Minkowski inequality, for any $x_{0}\in F$,
$\mathbb{E}_{\pi^{*}}[d^{4}(X,X_{t})]^{1/4}$ $\leq\mathbb{E}_{\pi^{*}}[d^{4}(x_{0},X_{t})]^{1/4}+\mathbb{E}_{\pi^{*}}[d^{4}(X,x_{0})]^{1/4}$
$\leq\mathbb{E}_{\pi}[d^{4}(x_{0},X_{t})]^{1/4}+\mathrm{diam}(F),$
where the last line follows since $\pi_{X_{t}}^{*}=\pi_{X_{t}}$, according to (3.26), and $d(x,x_{0})\leq\mathrm{diam}(F)$ for all $x\in F$. By (3.31), $\mathbb{E}_{\pi}[d^{4}(x_{0},X_{t})]\leq C^{4}+8tC^{2}.$ Thus,
$\mathbb{E}_{\pi^{*}}[d^{4}(X,X_{t})]^{1/4}\leq(C^{4}+8tC^{2})^{1/4}+\mathrm{diam}(F).$
Note that we have truncated the time $t$ up to the finite value $t_{0}$, implying $t\leq t_{0}$. Hence, the first variance requirement is satisfied. This gives the reason why we truncate $t$.
We next consider the second variance requirement. Substituting $x=x_{i}$ into (3.29) and taking summation over $i$, we obtain that for $\mu_{\mathbf{X}}$-a.e. $\mathbf{x}$,
$\sum_{i=1}^{n_{1}}\mathrm{Var}_{\mu}(d^{2}(x_{i},X_{i,t})|X_{i}=x_{i})\leq 8t\sum_{i=1}^{n_{1}}\mathbb{E}_{\mu}[d^{2}(x_{i},X_{i,t})|X_{i}=x_{i}]\leq 8tn(\tau-\delta),$
where the last inequality above follows by the expectation condition in (3.23). Since $t\leq t_{0}$, the second variance requirement is satisfied.
Since all requirements are satisfied by the transition probability induced by heat flow, we obtain a feasible solution to (3.22).
3.5.5. Step 5: Final Bound
We next analyze the relative entropy induced by the heat flow. Let
$D(t):=n_{0}D(\pi_{X_{t}|W}\|\nu|\pi_{W})+D(\mu_{\mathbf{X}_{t}}\|\nu^{\otimes n_{1}}).$
We first show that $D(t)\leq n\delta$ if $t=t_{0}$. Since $\nu$ admits a log-Sobolev inequality with constant $K$, by exponential decay in entropy [10, Theorem 5.2.1], we have that in this case,
(3.32) $D(t_{0})\leq e^{-2Kt_{0}}D(0)\leq e^{-2Kt_{0}}n\alpha=n\delta,$
where the last equality is due to the choice of $t_{0}$.
We next upper bound the relative entropy for $t=t_{1}$. By the chain rule for relative entropy and Fisher information and the fact that conditioning increases relative entropy, the following kind of tensorization property can be proven.
Theorem 9.
[59] Assume that $\nu$ admits a $\theta$-nonlinear log-Sobolev inequality for some convex function $\theta$. Then, $\nu^{\otimes n}$ admits a $\theta_{n}$-nonlinear log-Sobolev inequality where $\theta_{n}(t)=n\theta(t/n),\forall t\geq 0$.
By this theorem and Theorem 6, we obtain that
(3.33) $\sqrt{n\tau^{\prime\prime}}=\mathsf{W}_{2}(t)$ $\leq\int_{D(t)}^{D(0)}\frac{1}{\sqrt{n\theta(r/n)}}{\text{\rm d}}r=\sqrt{n}\int_{D(t)/n}^{D(0)/n}\frac{1}{\sqrt{\theta(t)}}{\text{\rm d}}t=\sqrt{n}\Upsilon\left(D(t)/n,\alpha\right).$
@@ 3.5.5. Step 5: Final Bound
Substituting (3.32) and (3.33) into (3.22) and letting $\delta\to 0$ yield that $\limsup_{n\to\infty}E_{n}(\alpha,\tau)\leq\bar{\psi}(\alpha,\tau).$